Fermi National Laboratory


Accelerator Update

Wednesday October 15 through Friday October 17
All shifts began in shutdown.

On Wednesday morning Linac was given permission to run beam. By late in the afternoon, the water group made the Main Injector (MI) LCW system operational and then began working on the Tevatron (TeV) LCW system. Early on the evening shift, Linac experts established beam to the 400 MeV area.

On Thursday morning, technicians began turning on the Booster RF. By late afternoon they had all eighteen stations up and running at approximately 60% voltage. Early on the evening shift, experts established beam to the Linac dump. Linac experts also reported that the newly installed Lambertson has a low enough field in the field free region between its on and its off state that no correction is needed.

During the Friday midnight shift, a few Booster RF stations tripped off, but operators turned them back on without incident.

RECYCLER
The Recycler ring bake outs are 68% complete (19 of 28). The gradient and quadrupole magnet heater tape repairs have finished. Techs have installed 91% of the additional vacuum readbacks (10 of 11). The four LEP correction magnets replacement and relocation have been completed. All cable pulls have been completed. Only about 5% of the Recycler alignment work has been completed. Rework of the 620 region is about 25% done.

TEVATRON
The TD crew continues to work on the B-3 conning tower replacement. They should finish this job by Monday.

Techs have completed the A2 magnet unrolls. The alignment group will check the twelve magnets with new stands.

Magnet bake outs will start on Friday or Saturday. Techs completed the installation of the new A-48 bypass. The bypass has been vacuum certified and interfacing continues. All the Japanese Pots vacuum tanks have been installed. The collimator will go in this week. Experts discovered a dipole magnet (47-5) that had a bad single phase, which is part of its cryo system, and had to be replaced.

The A-4 cryo system house is under going vacuum tests and certification.

All Lambertsons are baking.

The laser tracker crews have completed the TeV Net survey in sectors A, B, C, D, and F; a small part of E sector remains to be done.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/17/2003   email Fermilab